Presented below is the income statement of Total Inc. The income statement is based on sales of 100,000 units at $20 per unit.

Total estimates that 80% of cost of goods sold is variable, and 80% of operating expenses is fixed.

Sales $2,000,000
Less: COGS $600,000
Gross Margin $1,400,000
Less: Operating expenses $500,000
Net income $900,000

Total is considering lowering the sales price in order to increase sales. Management believes that if it reduces the selling price by 10%, then sales in units will increase by 10%.

Assuming Total reduces the selling price by 10%. How many units would Total have to sell to earn the same total contribution margin prior to the sell price reduction?
Answer #1

rate positively ..

Fixed Variable Per unit variable cost
Sales 2000000
COGS 600000 120000 480000          4.80
Operating expenses 500000 400000 100000          1.00
Net income 900000          5.80
Sales = 2000000
Variable cost = 580000
contribution margin 1420000
If sales price is reduced by 10%. Therefore revised sales price per unit = 20*90%= 18
Revised contribution margin = (18-5.8)/18 67.78%
Existing fixed cost = 520000
Contribution margin = 1420000
Total 1940000
Revised sales to achieve the contribution margin target = 1940000/67.78% 2862295
Price per unit = 18
Number of unit to sale = 159016